# A faster FPRAS for #NFA

### 20 Dec 2023 v1

@@ 1.1. Overview of our FPRAS
We remark that $\mathcal{L}({p}^{\ell-1})$ may not be disjoint for different predecessor states $p$. Hence, we cannot estimate $\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ by merely summing the individual estimates $\{\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})\,|\,(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta\}$. To address this, we maintain a set, $\mathsf{S}({p}^{\ell-1})$, for each $p$, which contains samples of strings from $\mathcal{L}({p}^{\ell-1})$. One possible idea is to construct the set $\mathsf{S}({p}^{\ell-1})$ using strings $w$ that have been sampled uniformly at random from $\mathcal{L}({p}^{\ell-1})$ but such a process introduces dependence among the sets $\mathsf{S}({p}^{\ell-1})$ for different $p$. The crucial observation in (Arenas et al., 2021) is that for any $k$, the following is true: for any word $w$ of length $k$, all words from $\mathcal{L}({q}^{\ell})$ with the suffix $w$ can be expressed as $\big{(}\bigcup\limits_{p\in P}\mathcal{L}({p}^{\ell-k})\big{)}\cdot\{w\}$ for some set $P$ of states, where $P$ depends only on $w,q,$ and $k$. This property is referred to as the self-reducible union property. This insight allows us to utilize the self-reducibility property for sampling. We can partition $\mathcal{L}(p^{\ell-1})$ into sets of strings where the last character is either ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}$ or ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}$, and compute estimates for each partition since each partition can be represented as a union of sets for which we (inductively) have access to both samples and estimates. We then sample the final character in proportion to the size of these estimates. In a similar recursive manner, we can sample strings, character by character. That is, we can sample strings by progressively extending suffixes of $w$ backwards! Of course, we must also account for the errors that propagate in the estimates of $\mathsf{N}({q}^{\ell})$.
@@
Our algorithm also relies on the self-reducible union property, and therefore, the basic structure of our algorithm is similar to that of (Arenas et al., 2021) on high-level. There are, however, significant and crucial technical differences that allow us to obtain an FPRAS with significantly lower time complexity. We highlight these differences below. As a first step, we highlight the two key invariants underpinning the algorithm and ensuing analysis in (Arenas et al., 2021):
  1. (1)
@@
    The following condition (Inv-ACJR) holds for every level $\ell$ (here, $\kappa=\frac{nm}{\epsilon}$), with high probability.
@@
    (Inv-ACJR) $\begin{array}[]{l}\xi(\ell)\triangleq\forall q{\in}Q,\forall P{\subseteq}Q,\big{|}\frac{|\mathsf{S}({q}^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathsf{S}({q}^{\ell})|}-\frac{|\mathcal{L}(q^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}\big{|}\leq\frac{1}{\kappa^{3}}\end{array}$
  2. (2)
@@
    $\mathsf{S}({q}^{\ell})$ is close in $(L_{\infty}$ distance)111 In (Arenas et al., 2021), the invariant is stated in a different, but equivalent, formulation: conditioned on $\xi(\ell)$, the set $\mathsf{S}({q}^{\ell})$ follows the distribution of a multi-set obtained by independently sampling every element of $\mathcal{L}(q^{\ell})$ uniformly at random, with replacement. to the distribution of multi-set obtained by independently sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
@@
  2. (2)
    The distribution of $\mathsf{S}({q}^{\ell})$ is close, in total variational distance, to the distribution of a multi-set constructed by sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
@@
For a given state $q$ and subset $P$, achieving a $\frac{1}{\kappa^{3}}$-additive approximation for $\frac{|\mathcal{L}(q^{\ell})\setminus\cup_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}$ with a probability of $1-\eta$ requires, according to standard concentration bounds, that the size of $\mathsf{S}({q}^{\ell})$ be $O(\kappa^{6}\log\eta^{-1})$. To ensure this additive approximation for all states $q$ and subsets $P$, one must consider union bounds over exponentially many events. Specifically, for the analysis to be successful, one must have $\eta^{-1}\in O(2^{mn})$. Consequently, the FPRAS of (Arenas et al., 2021) maintains $O(\kappa^{7})$ samples for each state $q$. In contrast, for $\texttt{AccurateN}_{q,\ell}$ to hold with a probability of $1-\eta$, our FPRAS only maintains $O\left(\frac{n^{4}}{\epsilon^{2}}\right)\log\eta^{-1}$ samples in $\mathsf{S}({q}^{\ell})$. Further, a union bound over only $O(mn)$ events is required, thus setting $\eta^{-1}=O(mn)$ is adequate.
The time complexity of an algorithm that follows the template outlined in Algorithm 1 (including both the FPRAS of (Arenas et al., 2021) as well as the FPRAS developed in this paper), can be expressed as $O(n\cdot(m\cdot n)\cdot(m^{2}n)\cdot|\mathsf{S}({q}^{\ell})|^{2})$ where the first $n$ from the length of every string, the factor $(m\cdot n)$ corresponds to the upper bound on the number of states in the unrolled automaton, and the factor $(m^{2}n)$ corresponds to the time taken for membership check. Accordingly, it is instructive to compare the bound of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ for $|S(q^{\ell})|$ in  (Arenas et al., 2021) to $O(\frac{n^{4}}{\epsilon^{2}})$ in our approach. Remarkably, the number of samples for every state in our approach is independent of $m$, the size of the automaton.
@@ 2. Preliminaries
We consider the binary alphabet $\Sigma=\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$. All results in this paper are extendable to alphabets of arbitrary but fixed constant size. A string $w$ over $\Sigma$ is either the empty string $\lambda$ (of length 00) or a non-empty finite sequence $w_{1}w_{2}\ldots w_{k}$ (of length $|w|=k>0$) where each $w_{i}\in\Sigma$. The set of all strings over $\Sigma$ is denoted by $\Sigma^{*}$. A non-deterministic finite automaton is a tuple $\mathcal{A}=(Q,I,\Delta,F)$ where $Q$ is a finite set of states, $I\in Q$ is the initial state, $\Delta\subseteq Q\times\Sigma\times Q$ is the transition relation, and $F$ is the set of accepting states. A run of $w$ on $\mathcal{A}$ is a sequence $\rho=q_{0},q_{1}\ldots,q_{k}$ such that $k=|w|$, $q_{0}=I$ and for every $i<k$, $(q_{i},w_{i+1},q_{i+1})\in\Delta$; $\rho$ is accepting if $q_{k}\in F$. The word $w$ is accepted by $\mathcal{A}$ if there is an accepting run of $w$ on $\mathcal{A}$. The set of strings accepted by $\mathcal{A}$ is $\mathcal{L}(\mathcal{A})$. For a natural number $n\in\mathbb{N}$, the $n^{\text{th}}$ slice of $\mathcal{A}$’s language, $\mathcal{L}(\mathcal{A}_{n})$, is the set of strings of length $n$ in $\mathcal{L}(\mathcal{A})$. For simplicity, this paper assumes a singleton set of accepting states, i.e., $F=\{q_{F}\}$ for some $q_{F}\in Q$. We introduce additional notation for our algorithms. For $q\in Q$ and $b\in\Sigma$, the $b$-predecessors of $q$ are:
$\textsf{Pred}(q,b)=\{p\,|\,(p,b,q)\in\Delta\}.$
For each $q\in Q$ and $1\leq\ell\leq n$, $\mathcal{L}(q^{\ell})$ denotes the set of strings $w$ of length $\ell$ with a run on $\mathcal{A}$ ending in $q$. The rest of the paper focuses on the automaton $\mathcal{A}=(Q,I,\Delta,F)$, and a number $n$ in unary, using $m$ to denote $|Q|$.
@@ 3. A Faster FPRAS
Given automaton $\mathcal{A}$ and a number $n$ in unary, we construct the unrolled automaton $\mathcal{A}_{\textsf{unroll}}$. We use the notation ${q}^{\ell}$ to denote the $\ell^{\text{th}}$ copy of state $q\in Q$. We use $\mathcal{L}({q}^{\ell})$ to denote the the set of all distinct strings $w$ of length $\ell$ for which there is a path (labeled by $w$) from the starting state to $q$ (i.e., those strings that go to state ${q}^{\ell}$ in $\mathcal{A}_{\textsf{unroll}}$). We will use $\mathsf{N}({q}^{\ell})$ to denote an estimate of $|\mathcal{L}({q}^{\ell})|$.
In Section 3.1, we formally present our FPRAS for #NFA, together with the statement of its correctness and running time (Theorem 3). Our FPRAS Algorithm 4 calls two subroutines Algorithm 3 and Algorithm 2, each of which are presented next, together with their informal descriptions. In later sections, we present the formal proofs of their correctness.
3.1. Description of the Algorithms
Algorithm 4 presents the main procedure for approximating the size of $\mathcal{L}(\mathcal{A})_{n}$, where $\mathcal{A}$ is the NFA and $n$ is the parameter that is given as input. The algorithm works in a dynamic programming fashion and maintains, for each state $q$ and each length $1\leq\ell\leq n$, a sampled subset $\mathsf{S}({q}^{\ell})$ of $\mathcal{L}({q}^{\ell}))$. As a key step, it uses the sampled sets corresponding to $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}})$ and $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}})$ to construct the desired sampled subset of $\mathcal{L}({q}^{\ell}))$. We first describe the two subroutines the $\texttt{AppUnion}{()}$ and $\mathsf{sample}()$. We start with describing the subroutines first.

### 08 Apr 2024 v2

@@ 1.1. Overview of our FPRAS
We remark that the sets $\{\mathcal{L}({p}^{\ell-1})\}_{p}$ may not be disjoint for different predecessor states $p$. Hence, we cannot estimate $\bigcup_{(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta}\mathcal{L}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})$ by merely summing the individual estimates $\{\mathsf{N}({p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}}^{\ell-1})\,|\,(p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta\}$. To address this, we maintain a set, $\mathsf{S}({p}^{\ell-1})$, for each $p$, which contains samples of strings from $\mathcal{L}({p}^{\ell-1})$. One possible idea is to construct the set $\mathsf{S}({p}^{\ell-1})$ using strings $w$ that have been sampled uniformly at random from $\mathcal{L}({p}^{\ell-1})$ but such a process introduces dependence among the sets $\mathsf{S}({p}^{\ell-1})$ for different $p$. The crucial observation in (Arenas et al., 2021) is that for any $k$, the following is true: for any word $w$ of length $k$, all words from $\mathcal{L}({q}^{\ell})$ with the suffix $w$ can be expressed as $\big{(}\bigcup\limits_{p\in P}\mathcal{L}({p}^{\ell-k})\big{)}\cdot\{w\}$ for some set $P$ of states, where $P$ depends only on $w,q,$ and $k$. This property is referred to as the self-reducible union property. This insight allows us to utilize the self-reducibility property for sampling. We can partition $\mathcal{L}(p^{\ell-1})$ into sets of strings where the last character is either ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}$ or ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}$, and compute estimates for each partition since each partition can be represented as a union of sets for which we (inductively) have access to both samples and estimates. We then sample the final character in proportion to the size of these estimates. Similarly, we can sample strings, character by character. That is, we can sample strings by progressively extending suffixes of $w$ backwards! Of course, we must also account for the errors that propagate in the estimates of $\mathsf{N}({q}^{\ell})$.
@@
Since our algorithm also relies on the self-reducible union property, the basic structure of our algorithm is similar to that of (Arenas et al., 2021) on high-level. There are, however, significant and crucial technical differences that allow us to obtain an FPRAS with significantly lower time complexity. We highlight these differences below by first highlighting the two key properties underpinning the algorithm and ensuing analysis in (Arenas et al., 2021):
  1. (ACJR-1).
@@
    For every level $\ell$, the following condition holds with high probability (here, $\kappa=\frac{nm}{\epsilon}$):
@@
    $\begin{array}[]{l}\xi(\ell)\triangleq\forall q{\in}Q,\forall P{\subseteq}Q,\ \left|\frac{|\mathsf{S}({q}^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathsf{S}({q}^{\ell})|}-\frac{|\mathcal{L}(q^{\ell})\setminus\bigcup\limits_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}\right|\leq\frac{1}{\kappa^{3}}\end{array}$
  2. (ACJR-2).
@@
    $\mathsf{S}({q}^{\ell})$ is close in $(L_{\infty}$ distance)222 In (Arenas et al., 2021), the invariant is stated in a different, but equivalent, formulation: conditioned on $\xi(\ell)$, the set $\mathsf{S}({q}^{\ell})$ follows the distribution of a multi-set obtained by independently sampling every element of $\mathcal{L}(q^{\ell})$ uniformly at random, with replacement. to the distribution of multi-set obtained by independently sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
@@
  2. (Inv-2).
    The distribution of $\mathsf{S}({q}^{\ell})$ is close, in total variation distance, to the distribution of a multi-set constructed by sampling (with replacement) every element of $\mathcal{L}(q^{\ell})$ uniformly at random.
@@
For a given state $q$ and subset $P$, achieving a $\frac{1}{\kappa^{3}}$-additive approximation for $\frac{|\mathcal{L}(q^{\ell})\setminus\cup_{p\in P}\mathcal{L}(p^{\ell})|}{|\mathcal{L}(q^{\ell})|}$ with a probability of $1-\eta$ requires, according to standard concentration bounds, that the size of $\mathsf{S}({q}^{\ell})$ be $O(\kappa^{6}\log\eta^{-1})$. To ensure this additive approximation for all states $q$ and subsets $P$, one must consider union bounds over exponentially many events. Specifically, for the analysis to be successful, one must have $\eta^{-1}\in O(2^{mn})$. Consequently, the FPRAS of (Arenas et al., 2021) maintains $O(\kappa^{7})$ samples for each state $q$. In contrast, for $\texttt{AccurateN}_{q,\ell}$ to hold with a probability of $1-\eta$, our FPRAS only maintains $O\left(\frac{n^{4}}{\epsilon^{2}}\right)\log\eta^{-1}$ samples in $\mathsf{S}({q}^{\ell})$. Furthermore, a union bound over only $O(mn)$ events is required, thus setting $\eta^{-1}=O(mn)$ is adequate.
The time complexity of an algorithm that follows the template outlined in Fig. 1 (including both the FPRAS of (Arenas et al., 2021) as well as the FPRAS developed in this paper), has a quadratic depends on $|\mathsf{S}({q}^{\ell})|$. Accordingly, it is instructive to compare the bound of $O(\frac{m^{7}n^{7}}{\epsilon^{7}})$ for $|S(q^{\ell})|$ in  (Arenas et al., 2021) to $O(\frac{n^{4}}{\epsilon^{2}})$ in our approach. Remarkably, the number of samples for every state in our approach is independent of $m$, the size of the automaton.
Organization. In Section 2 we recall some helpful background as well as introduce a key concept (distribution entanglement) useful for the rest of the paper. In Section 3, we describe our algorithm, together with two auxiliary algorithms and state their formal correctness guarantees, whose analysis we present in Section 4. We conclude in Section 5.
@@ 2. Preliminaries
NFA and words. We consider the binary alphabet $\Sigma=\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}$. All results in this paper are extendable to alphabets of arbitrary but fixed constant size. A string $w$ over $\Sigma$ is either the empty string $\lambda$ (of length 00) or a non-empty finite sequence $w_{1}w_{2}\ldots w_{k}$ (of length $|w|=k>0$) where each $w_{i}\in\Sigma$. The set of all strings over $\Sigma$ is denoted by $\Sigma^{*}$. A non-deterministic finite automaton is a tuple $\mathcal{A}=(Q,I,\Delta,F)$ where $Q$ is a finite set of states, $I\in Q$ is the initial state, $\Delta\subseteq Q\times\Sigma\times Q$ is the transition relation, and $F$ is the set of accepting states. A run of $w$ on $\mathcal{A}$ is a sequence $\rho=q_{0},q_{1}\ldots,q_{k}$ such that $k=|w|$, $q_{0}=I$ and for every $i<k$, $(q_{i},w_{i+1},q_{i+1})\in\Delta$; $\rho$ is accepting if $q_{k}\in F$. The word $w$ is accepted by $\mathcal{A}$ if there is an accepting run of $w$ on $\mathcal{A}$. The set of strings accepted by $\mathcal{A}$ is $\mathcal{L}(\mathcal{A})$. For a natural number $n\in\mathbb{N}$, the $n^{\text{th}}$ slice of $\mathcal{A}$’s language, $\mathcal{L}(\mathcal{A}_{n})$, is the set of strings of length $n$ in $\mathcal{L}(\mathcal{A})$. For simplicity, this paper assumes a singleton set of accepting states, i.e., $F=\{q_{F}\}$ for some $q_{F}\in Q$. For $q\in Q$ and ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\Sigma$, the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessors of $q$ are given by the set $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}})=\{p\,|\,(p,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}},q)\in\Delta\}.$ Given automaton $\mathcal{A}$ and a number $n$ in unary, we construct the unrolled automaton $\mathcal{A}_{\textsf{unroll}}$, and assume that all states in $\mathcal{A}_{\textsf{unroll}}$ are reachable from the initial state. We use the notation ${q}^{\ell}$ to denote the $\ell^{\text{th}}$ copy of state $q\in Q$. For each state $q$ and level $0\leq\ell\leq n$, we use $\mathcal{L}({q}^{\ell})$ to denote the the set of all distinct strings $w$ of length $\ell$ for which there is a path (labeled by $w$) from the starting state to $q$.
Total Variation Distance.The total variation distance between two random variables $X,Y$ over domain $\Omega$ is defined as:
$\mathsf{TV}(X,Y)=\sum_{\omega\in\Omega}\Pr[X=\omega]-\min(\Pr[X=\omega],\Pr[Y=\omega]).$
2.1. A New Notation: Distribution Entanglement
We introduce a new notation, referred to as distribution entanglement, to argue about the probability of events for cases when the distribution of a certain random variable were to mimic distribution of another random variable. For an event $\mathcal{E}$ and random variables $X$ and $Y$, we use ${\mathsf{Pr}}_{X}\!\left[\mathcal{E};{Y}\right]$ to denote the probability of $\mathcal{E}$ when $X$ has distribution entangled to $Y$, which is defined as follows:
${\mathsf{Pr}}_{X}\!\left[\mathcal{E};{Y}\right]=\sum_{\omega\in\Omega}{\mathsf{Pr}}\!\left[Y=\omega\right]\cdot{\mathsf{Pr}}\!\left[\mathcal{E}|X=\omega\right].$
where $\Omega$ is the support of $Y$. The above notion is well-defined only when it is the case for all $\omega\in\Omega$, we have $\Pr[X=\omega]>0$. Likewise, we can also extend this notion for conditioned events. That is, for events $\mathcal{E}$, $\mathcal{F}$ and random variables $X$ and $Y$, the conditional probability of $\mathcal{E}$ given $\mathcal{F}$ when $X$ has distribution entangled to $Y$is defined as follows:
${\mathsf{Pr}}_{X}\!\left[\mathcal{E}|\mathcal{F};{Y}\right]=\sum_{\omega\in\Omega}{\mathsf{Pr}}\!\left[Y=\omega\right]\cdot{\mathsf{Pr}}\!\left[\mathcal{E}|\mathcal{F}\cap X=\omega\right].$
Whenever $X$ is clear from the context, we will omit it from the notation.
@@ 3. A Faster FPRAS
In this section we formally present our FPRAS for #NFA, together with the statement of its correctness and running time (Theorem 3). Our FPRAS Algorithm 3 calls two subroutines AppUnion (Algorithm 1) and sample (Algorithm 2), each of which are presented next, together with their informal descriptions and formal statements of correctness. Algorithm 3 presents the main procedure for approximating the size of $\mathcal{L}(\mathcal{A}_{n})$, where $\mathcal{A}$ is the NFA and $n$ is the parameter that is given as input. The algorithm works in a dynamic programming fashion and maintains, for each state $q$ and each length $1\leq\ell\leq n$, a sampled subset $\mathsf{S}({q}^{\ell})$ of $\mathcal{L}({q}^{\ell})$. As a key step, it uses the sampled sets corresponding to $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}})$ and $\textsf{Pred}(q,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}})$ to construct the desired sampled subset of $\mathcal{L}({q}^{\ell})$. We first describe the two subroutines the AppUnion and sample.
3.1. Approximating Union of Sets
AppUnion (Algorithm 1) takes parameters $\epsilon,\delta$ and inputs $\epsilon_{\textsf{sz}}$ and (access to) sets $T_{1},\dots,T_{k}$. For each set $T_{i}$ the access is given in the form of a membership oracle $O_{i}$, a subset $S_{i}$ (obtained by sampling $T_{i}$ with replacements) presented as a list, and an estimate $\textsf{sz}_{i}$ of the size of $T_{i}$. Using these, Algorithm 1 outputs an $(\epsilon,\delta)$ estimate of the size of the union $\cup_{i=1}^{k}T_{i}$.
The precise algorithm presented here represents a modification of the classic Monte Carlo-based scheme due to Karp and Luby (Karp and Luby, 1985). The proof of its correctness also shares similarities with (Karp and Luby, 1985) and we summarize the intuition here. For a $\sigma\in\cup_{i=1}^{k}T_{i}$, let $L(\sigma)$ denote the smallest index $i\in\{1,2,\ldots,k\}$ such that $\sigma\in T_{i}$. Now, let $\mathcal{U}_{\textsf{unique}}$ be the set of all pairs $(\sigma,L(\sigma))$ and let $\mathcal{U}_{\textsf{multiple}}$ be the set of all $(\sigma,i)$ pairs such that $\sigma\in T_{i}$. Consequently, the cardinality of the set $\mathcal{U}_{\textsf{unique}}$ is $|\cup_{i=1}^{k}T_{i}|$, while the cardinality of $\mathcal{U}_{\textsf{multiple}}$ is $\sum_{i=1}^{k}|T_{i}|$. By drawing sufficiently-many ($t$, to be precise) samples from $\mathcal{U}_{\textsf{multiple}}$ and assessing the fraction belonging to $\mathcal{U}_{\textsf{unique}}$, one can estimate $|\mathcal{U}_{\textsf{unique}}|/|\mathcal{U}_{\textsf{multiple}}|$, which when multiplied by $|\mathcal{U}_{\textsf{multiple}}|$ provides an estimate for the size of $\mathcal{U}_{\textsf{unique}}$, which is the desired output.